# LLM-Select: Feature Selection with Large Language Models

### v1

@@ A.2 Feature Selection, Model Training, and Hyperparameter Optimization
As described in Section 4.1 of the main text, we evaluate the effectiveness of each feature selection method by measuring how the test performance of a downstream -penalized logistic regression (for classification tasks) or linear regression (for regression tasks) model changes as we vary the proportion of features selected from 10% to 100%, in approximately 10% increments. For logistic regression, we minimize the negative log-likelihood using the L-BFGS optimizer (Zhu et al., 1997) and use importance weighting to balance the weights of the positive and negative samples, as most of the datasets exhibit label imbalance. For linear regression, we minimize the mean squared error (MSE) using the L-BFGS optimizer. For all experiments, we aggregate the results over 5 random seeds (=[1,2,3,4,5]), which control the train-validation splits and the behavior of random feature selection.
@@ A.2.1 Additional Details on Feature Selection Methods
For feature selection with the LASSO (Tibshirani, 1996) on each dataset, we first train an -penalized logistic regression model with inverse regularization coefficient (for classification tasks) or an -penalized linear regression model with regularization coefficient (for regression tasks)181818For logistic regression, we configure the inverse regularization coefficient instead of the regularization coefficient , as the scikit-learn implementation of logistic regression uses the former.. To compute the sparse-to-dense regularization paths with warm starts (Friedman et al., 2010), we then gradually increase by iteratively multiplying powers of 1.02 (i.e., ) and gradually decrease by iteratively multiplying powers of (i.e., ) until all features become active. For logistic regression, we train each model by minimizing the negative log-likelihood using the SAGA optimizer (Defazio et al., 2014). For linear regression, we train each model by minimizing the MSE using the SAGA optimizer. After computing the full regularization path, we identify the regularization strengths that select 10% to 100% of all features in approximately 10% increments.
@@ A.3.1 Comparison of LLM-Score, LLM-Rank, and LLM-Seq to Baselines
In Figure A2, we observe that LLM-Score, LLM-Rank, and LLM-Seq all achieve strong feature selection performance competitive with the data-driven baselines. For the larger models, notably GPT-4 (in red) and GPT-3.5 (in green), we consistently observe such strong performance across many datasets. For the smaller models, notably the Llama-2 models (in blue, orange, and purple), we observe that feature selection performance is generally more sensitive to the choice of selection strategy (i.e., LLM-Rank or LLM-Seq), but that performance can be strong when an adequate selection strategy is used. For example, on the Credit-G and COMPAS Recidivism datasets, we see that using LLM-Rank significantly improves the performances of GPT-3.5 and the Llama-2 models over that achieved with LLM-Score (shown in Figure 2)—from being close to the average performance of random feature selection (in black) to being on par or better than those of the best-performing baselines. For LLM-Seq, we observe a similar improvement on the Pima Indians Diabetes and COMPAS Recidivism datasets.
Table A1 further illustrates these findings from Figure A2. We see that for the larger GPT-4 and GPT-3.5 models, the area of the best-performing LLM-based feature selection method is often on par with the area of the best-performing baseline, across most datasets. For example, on the Credit-G dataset, we see that the best-performing baseline (MRMR) achieves an area of 0.7496, while the best-performing selection strategy for GPT-4 (LLM-Seq) outperforms it with an area of 0.7710, and the best-performing selection strategy for GPT-3.5 (LLM-Rank) performs as well with an area of 0.7495. We observe that these models achieve strong performance on the other datasets as well (e.g., COMPAS Recidivism, Pima Indians Diabetes, Wine Quality). Meanwhile, for the smaller Llama-2 models, we see that the area of the best-performing selection strategy can be on par with that of the best-performing baseline but that feature selection performance more sensitive to the choice of selection strategy. For example, on the Pima Indians Diabetes dataset, we see that the areas of the best-performing selection strategies for the three Llama-2 models are 0.7728 (70B), 0.7744 (13B), and 0.7709 (7B), respectively, which are on par with the area of 0.7855 for the best-performing baseline (RFE). However, as noted in Result 1, Section 4.1 of the main text, the performance of the smaller models are less consistent across datasets and the 3 LLM-based feature selection methods.
@@ A.3.2 Comparison of LLM-Score, LLM-Rank, and LLM-Seq Across LLMs of Varying Scale
In this section, we include additional results that compare the feature selection behaviors and performances of LLM-Score, LLM-Rank, and LLM-Seq across LLMs of varying scale. We note that all of the results shown in this section are based on the LLM outputs generated from the default prompt, which does not include any dataset-specific context or few-shot examples, and with greedy decoding. As done for GPT-4 in Result 2, Section 4.1 of the main text (Figure 3), we show the changes in average test performance for LLM-Score, LLM-Seq, and LLM-Rank for GPT-3.5 and the Llama-2 models in Figures A3–A6. In Figure A8, we also show the average rank correlation (Kendall’s ) between (i) LLM-Score and LLM-Rank (left) and that between (ii) LLM-Score and LLM-Seq (right) to show how well the orderings of the features by their relevance to the prediction target agree between each pair selection strategies being compared. As described in Result 4, Section 4.1 of the main text, the Kendall’s coefficient quantifies the agreement in pairwise orderings, where +1/-1 indicates perfect agreement/disagreement.
Together with Figure 3 discussed in Result 2, Section 4.1 of the main text, Figures A3–A6 show that the feature selection paths of LLM-Score, LLM-Rank, and LLM-Seq are generally more consistent with one another (i.e., exhibits less variability) with increasing model scale, with GPT-4 showing almost perfect alignment on all datasets except Diabetes Progression. Figure A8 also shows that the average rank correlation between LLM-Score and LLM-Rank and that between LLM-Score and LLM-Seq both tend to increase with increasing model scale. Table A1 further illustrates this finding, where we see that with increasing model scale, the areas under the feature selection paths for LLM-Score, LLM-Rank, and LLM-Seq tend to be more similar.
@@ A.3.5 Feature Subset Initialization for LLM-Seq
In this section, we provide additional results that compare the performances of LLM-Seq when (i) starting with an empty feature subset or (ii) starting with a feature subset containing the highest scoring feature according to LLM-Score. See Section 3 of the main text for a detailed description of the difference between the two settings. In Figures A9–A13, we show the changes in the downstream average test performance for LLM-Seq based on GPT-4, GPT-3.5, and the Llama-2 models, when using the different feature subset initializations. The results show that with the exception of the largest GPT-4 model (Figure A9), the choice of feature subset initialization generally has a noticeable impact on the feature selection performance of LLM-Seq. In Table A3, we report the areas under the feature selection paths shown in Figures A9–A13 to quantify and compare the feature selection performances for the two initializations. For each dataset and each LLM, the results in Table A3 show that starting with an empty feature subset often leads to better feature selection performance.
@@
A.3.6 Investigation of the Semantics of LLM-Score
@@ B.1 Datasets
We construct supersets of the original Income, Employment, Public Coverage, and Mobility binary classification datasets from folktables (Ding et al., 2021) by extracting all available features available from the 2018 American Community Survey (ACS) Public Use Microdata Sample (PUMS) data191919https://www.census.gov/programs-surveys/acs/microdata.html for California households. The full ACS PUMS dataset contains a total of 286 features, from which we exclude (i) the feature that is used for defining the target label, (ii) features that serve as unique identifiers for each sample, and (iii) features that lead to label leakage when included. Below, we provide the remaining details for each folktables dataset:
@@ B.1.2 MIMIC-IV Datasets
First and only ICU stay: Following standard practice (Wang et al., 2020), if a patient has multiple ICU stays recorded in the database across all hospitalizations, we only consider the first-ever ICU stay. This ensures that there are no ICU stays in Additionally, as ICD diagnosis codes are assigned to each hospitalization, which may include multiple ICU stays, we only include first ICU stays that do not have subsequent ICU stays recorded within the same hospitalization to ensure that the ICD diagnosis codes correspond to them.
@@ B.2 Feature Selection, Model Training, and Hyperparameter Optimization
As described in Section 4.2 of the main text, we evaluate the effectiveness of each feature selection method by measuring the test performance of a downstream prediction model when selecting only 30% of all input concepts. For downstream training, we minimize the binary cross-entropy loss using the Adam optimizer (Kingma & Ba, 2015), with the default momentum hyperparameters , and . For all datasets, we use importance weighting together with early stopping when training the downstream model to balance the weights of the positive and negative samples, as most of the folktables and MIMIC-IV datasets exhibit label imbalance. For all experiments, we aggregate the results over 5 random seeds (=[1,2,3,4,5]), which control the train-validation splits, the hyperparameter samples generated for random-search hyperparameter optimization for the downstream model, and the initialization of the model parameters.
@@ B.2.1 Additional Details on Feature Selection Methods
gLASSO.
For feature selection with group LASSO (gLASSO; Yuan & Lin, 2006), we first train an unregularized logistic regression model on each dataset. To compute the sparse-to-dense regularization paths with warm starts (Friedman et al., 2010), we then gradually increase the regularization coefficient as [0.001, 0.005, 0.01, 0.02, 0.03, 0.04, 0.05, 0.06, 0.07, 0.08, 0.09, 0.1, 0.15, 0.2, 0.25, 0.3, 0.35, 0.4, 0.45, 0.5], where for each value of , we train a new model initialized with the previously learned model parameters. After computing the regularization path, we identify the regularization strength that selects approximately the same number of features as LLM-Score on each dataset, and use the active features with nonzero weights for training the downstream prediction model. Figure B1(b) shows the gLASSO regularization paths computed for all of the datasets, along with the regularization strength used for feature selection (marked with a star).
@@ B.2.2 Downstream Model Training and Hyperparameter Optimization
After selecting the features to use for training according to each feature selection method, we run a random search with asynchronous successive halving (Li et al., 2020) to select the best hyperparameter configuration to use for training and evaluation of each downstream prediction model. For each downstream prediction model, we randomly sample 40 different hyperparameter configurations, and select the configuration that maximizes the validation AUROC. We then measure the performance of the selected model on the test set. We repeat this process 5 times with different random seeds (=[1,2,3,4,5]) and average the results. Below, we provide the hyperparameter search space used for each model. For any hyperparameter that is not explicitly listed below for LightGBM (Ke et al., 2017), we use the default value used in the corresponding API202020https://lightgbm.readthedocs.io/en/stable/Parameters.html.
@@ Appendix C Prompting for Feature Importance Scores
In practice, we found that choosing the right output format for a given LLM is also important to consistently obtain a well-structured (i.e., easy to parse and extract the relevant information) and adequate response for the LLM, especially for smaller models like Llama-2 with 7B parameters. We only consider the output formats available off-the-shelf from the LangChain API212121https://github.com/langchain-ai/langchain, which we used to implement most of our LLM prompting methods. Below, we provide the output format instructions used for the different LLMs:

### v2

@@ A.2 Feature Selection, Model Training, and Hyperparameter Optimization
As described in Section 4.1 of the main text, we evaluate the effectiveness of each feature selection method by measuring how the test performance of a downstream -penalized logistic regression (for classification tasks) or linear regression (for regression tasks) model changes as we vary the proportion of features selected from 10% to 100%, in approximately 10% increments. For logistic regression, we minimize the negative log-likelihood using the L-BFGS optimizer (Zhu et al., 1997) and use importance weighting to balance the weights of the positive and negative samples, as some of the datasets exhibit label imbalance. For linear regression, we minimize the mean squared error (MSE) using the L-BFGS optimizer. For all experiments, we aggregate the results over 5 random seeds (=[1,2,3,4,5]), which control the train-validation splits and the behavior of random feature selection, in order to ensure the robustness of results.
@@ A.2.1 Additional Details on Feature Selection Methods
For feature selection with the LASSO (Tibshirani, 1996) on each dataset, we first train an -penalized logistic regression model with inverse regularization coefficient (for classification tasks) or an -penalized linear regression model with regularization coefficient (for regression tasks)191919For logistic regression, we configure the inverse regularization coefficient instead of the regularization coefficient , as the scikit-learn implementation of logistic regression uses the former.. To compute the sparse-to-dense regularization paths with warm starts (Friedman et al., 2010), we then gradually increase by iteratively multiplying powers of 1.02 (i.e., ) and gradually decrease by iteratively multiplying powers of (i.e., ) until all features become active. For logistic regression, we train each model by minimizing the negative log-likelihood using the SAGA optimizer (Defazio et al., 2014). For linear regression, we train each model by minimizing the MSE using the SAGA optimizer. After computing the full regularization path, we identify the regularization strengths that select 10% to 100% of all features in approximately 10% increments.
@@ A.3.1 Comparison of LLM-Score, LLM-Rank, and LLM-Seq to Baselines
In Figure A2, we observe that LLM-Score, LLM-Rank, and LLM-Seq all achieve strong feature selection performance competitive with the data-driven baselines. For the larger models, notably GPT-4 (in blue) and GPT-3.5 (in green), we consistently observe such strong performance across many datasets. For the smaller models, notably the Llama-2 models (in orange, pink, and yellow), we observe that feature selection performance is generally more sensitive to the choice of selection strategy (i.e., LLM-Rank or LLM-Seq), but that performance can be strong when an adequate selection strategy is used. For example, on the Credit-G and COMPAS Recidivism datasets, we see that using LLM-Rank significantly improves the performances of GPT-3.5 and the Llama-2 models over that achieved with LLM-Score (shown in Figure 2)—from being close to the average performance of random feature selection (in black) to being on par or better than those of the best-performing baselines. For LLM-Seq, we observe a similar improvement on the Pima Indians Diabetes and COMPAS Recidivism datasets.
Table A1 further illustrates these findings from Figure A2. We see that for the larger GPT-4 and GPT-3.5 models, the area of the best-performing LLM-based feature selection method is often on par with the area of the best-performing baseline, across most datasets. For example, on the Credit-G dataset, we see that the best-performing baseline (MRMR) achieves an area of 0.7496, while the best-performing selection strategy for GPT-4 (LLM-Seq) outperforms it with an area of 0.7710, and the best-performing selection strategy for GPT-3.5 (LLM-Rank) performs as well with an area of 0.7495. We observe that these models achieve strong performance on the other datasets as well (e.g., COMPAS Recidivism, Pima Indians Diabetes, Wine Quality). Meanwhile, for the smaller Llama-2 models, we see that the area of the best-performing selection strategy can be on par with that of the best-performing baseline but that feature selection performance more sensitive to the choice of selection strategy. For example, on the Pima Indians Diabetes dataset, we see that the areas of the best-performing selection strategies for the three Llama-2 models are 0.7728 (70B), 0.7744 (13B), and 0.7709 (7B), respectively, which are on par with the area of 0.7855 for the best-performing baseline (RFE). However, as noted in Result 1, Section 4.1 of the main text, the performance of the smaller models are less consistent across datasets and the three LLM-based feature selection methods.
@@ A.3.2 Comparison of LLM-Score, LLM-Rank, and LLM-Seq Across LLMs of Varying Scale
In this section, we include additional results that compare the feature selection behaviors and performances of LLM-Score, LLM-Rank, and LLM-Seq across LLMs of varying scale. We note that all of the results shown in this section are based on the LLM outputs generated from the default prompt, which does not include any dataset-specific context or few-shot examples, and with greedy decoding. As done for GPT-4 and GPT-3.5 in Result 2, Section 4.1 of the main text (Figure 3), we show the changes in average test performance for LLM-Score, LLM-Seq, and LLM-Rank for GPT-3.5 and the Llama-2 models in Figures A3–A5. In Figure A7, we also show the average rank correlation (Kendall’s ) between (i) LLM-Score and LLM-Rank (left) and that between (ii) LLM-Score and LLM-Seq (right) to show how well the orderings of the features by their relevance to the prediction target agree between each pair selection strategies being compared. As described in Result 4, Section 4.1 of the main text, the Kendall’s coefficient quantifies the agreement in pairwise orderings, where +1/-1 indicates perfect agreement/disagreement.
Together with Figure 3 discussed in Result 2, Section 4.1 of the main text, Figures A3–A5 show that the feature selection paths of LLM-Score, LLM-Rank, and LLM-Seq are generally more consistent with one another (i.e., exhibits less variability) with increasing model scale, with GPT-4 showing almost perfect alignment on all datasets except Diabetes Progression. Figure A7 also shows that the average rank correlation between LLM-Score and LLM-Rank and that between LLM-Score and LLM-Seq both tend to increase with increasing model scale. Table A1 further illustrates this finding, where we see that with increasing model scale, the areas under the feature selection paths for LLM-Score, LLM-Rank, and LLM-Seq tend to be more similar.
@@ A.3.5 Feature Subset Initialization for LLM-Seq
In this section, we provide additional results that compare the performances of LLM-Seq when (i) starting with an empty feature subset or (ii) starting with a feature subset containing the highest scoring feature according to LLM-Score. See Section 3 of the main text for a detailed description of the difference between the two settings. In Figures A8–A12, we show the changes in the downstream average test performance for LLM-Seq based on GPT-4, GPT-3.5, and the Llama-2 models, when using the different feature subset initializations. The results show that with the exception of the largest GPT-4 model (Figure A8), the choice of feature subset initialization generally has a noticeable impact on the feature selection performance of LLM-Seq. In Table A3, we report the areas under the feature selection paths shown in Figures A8–A12 to quantify and compare the feature selection performances for the two initializations. For each dataset and each LLM, the results in Table A3 show that starting with an empty feature subset often leads to better feature selection performance.
@@
A.3.6 Importance of Text Semantics for LLM-based Feature Selection
In this section, we provide additional results that demonstrate the importance of text descriptions with rich real-world semantics (e.g., column names in tabular datasets, auxiliary metadata) for LLM-based feature selection. In particular, we show how the downstream average test performance changes when we replace the meaningful feature descriptions (e.g., “gender”, “height”) in each dataset with irrelevant names (e.g., “feature_1”, “feature_2”) and use GPT-4-based LLM-Score to select features. In Table A4, we report the areas under the feature selection paths to quantify and compare the feature performances for the two settings. We also show the results for the random feature selection baseline, as we would expect the behavior of LLM-Score to be similar to random selection in the absence of meaningful text descriptions. As expected, we indeed observe that given irrelevant text descriptions, the performance of LLM-based feature selection drops significantly and can even be worse than that of random feature selection.
| Classification Datasets | Regression Datasets | |||||||||||||
| Give Me | COMPAS | Pima | CA | Diabetes | Wine | Miami | NYC | |||||||
| GPT-4 (Desc.) | 0.7975 | 0.8593 | 0.7423 | 0.8305 | 0.7794 | 0.9149 | 0.8654 | 0.6278 | 44.3943 | 0.5790 | 0.2423 | 1.4121 | 0.0618 | 1.9690 |
| GPT-4 (No Desc.) | 0.5707 | 0.6129 | 0.7317 | 0.6224 | 0.7307 | 0.7190 | 0.6308 | 0.8955 | 51.2819 | 0.6447 | 0.3071 | 1.7848 | 1.6954 | 4.1610 |
A.3.7 Investigation of the Semantics of LLM-Score
@@ B.1 Datasets
We construct supersets of the original Income, Employment, Public Coverage, and Mobility binary classification datasets from folktables (Ding et al., 2021) by extracting all available features available from the 2018 American Community Survey (ACS) Public Use Microdata Sample (PUMS) data202020https://www.census.gov/programs-surveys/acs/microdata.html for California households. The full ACS PUMS dataset contains a total of 286 features, from which we exclude (i) the feature that is used for defining the target label, (ii) features that serve as unique identifiers for each sample, and (iii) features that lead to label leakage when included. Below, we provide the remaining details for each folktables dataset:
@@ B.1.2 MIMIC-IV Datasets
First and only ICU stay: Following standard practice (Wang et al., 2020), if a patient has multiple ICU stays recorded in the database across all hospitalizations, we only consider the first-ever ICU stay. Additionally, as ICD diagnosis codes are assigned to each hospitalization, which may include multiple ICU stays, we only include first ICU stays that do not have subsequent ICU stays recorded within the same hospitalization to ensure that the ICD diagnosis codes correspond to them.
@@ B.2 Feature Selection, Model Training, and Hyperparameter Optimization
As described in Section 4.2 of the main text, we evaluate the effectiveness of each feature selection method by measuring the test performance of a downstream prediction model when selecting only 30% of all input concepts. For downstream training, we minimize the binary cross-entropy loss using the Adam optimizer (Kingma & Ba, 2015) with the default hyperparameters , and . For all datasets, we use importance weighting together with early stopping when training the downstream model to balance the weights of the positive and negative samples, as most of the folktables and MIMIC-IV datasets exhibit label imbalance. For all experiments, we aggregate the results over 5 random seeds (=[1,2,3,4,5]), which control the train-validation splits, the hyperparameter samples generated for random-search hyperparameter optimization for the downstream model, and the initialization of the model parameters.
@@ B.2.1 Additional Details on Feature Selection Methods
LASSO.
For feature selection with the LASSO (Tibshirani, 1996) with group-wise feature sparsity (Yuan & Lin, 2006), we first train an unregularized logistic regression model on each dataset. To compute the sparse-to-dense regularization paths with warm starts (Friedman et al., 2010), we then gradually increase the regularization coefficient as [0.001, 0.005, 0.01, 0.02, 0.03, 0.04, 0.05, 0.06, 0.07, 0.08, 0.09, 0.1, 0.15, 0.2, 0.25, 0.3, 0.35, 0.4, 0.45, 0.5], where for each value of , we train a new model initialized with the previously learned model parameters. After computing the regularization path, we identify the regularization strength that selects approximately the same number of features as LLM-Score on each dataset, and use the active features with nonzero weights for training the downstream prediction model. Figure B1(b) shows the LASSO regularization paths computed for all of the datasets, along with the regularization strength used for feature selection (marked with a star).
@@ B.2.2 Downstream Model Training and Hyperparameter Optimization
After selecting the features to use for training according to each feature selection method, we run a random search with asynchronous successive halving (Li et al., 2020) to select the best hyperparameter configuration to use for training and evaluation of each downstream prediction model. For each downstream prediction model, we randomly sample 40 different hyperparameter configurations, and select the configuration that maximizes the validation AUROC. We then measure the performance of the selected model on the test set. We repeat this process 5 times with different random seeds (=[1,2,3,4,5]) and average the results. Below, we provide the hyperparameter search space used for each model. For any hyperparameter that is not explicitly listed below for LightGBM (Ke et al., 2017), we use the default value used in the corresponding API212121https://lightgbm.readthedocs.io/en/stable/Parameters.html.
@@ Appendix C Prompting for Feature Importance Scores
In practice, we found that choosing the right output format for a given LLM is also important to consistently obtain a well-structured (i.e., easy to parse and extract the relevant information) and adequate response for the LLM, especially for smaller models like Llama-2 with 7B parameters. We only consider the output formats available off-the-shelf from the LangChain API222222https://github.com/langchain-ai/langchain, which we used to implement most of our LLM prompting methods. Below, we provide the output format instructions used for the different LLMs: